OpenStudy (here_to_help15):

What is the domain of the function y=_/6x+42

OpenStudy (here_to_help15):

i meant y=3\[\sqrt{6x+42}\]

OpenStudy (here_to_help15):

3\[\sqrt{6x+42}\]

OpenStudy (here_to_help15):

3 goes before square

OpenStudy (here_to_help15):

@camerondoherty

OpenStudy (camerondoherty):

You mean: \[^3\sqrt{6x+42}\]

OpenStudy (here_to_help15):

yes

OpenStudy (camerondoherty):

The domain is all real numbers

OpenStudy (here_to_help15):

y did u give me medal?

OpenStudy (here_to_help15):

@camerondoherty look at your messages plz